# Computing hydration free energies of small molecules with first principles accuracy

### v2

@@ II.2 Hydration free energy calculations
II.2.1 Organic molecule benchmark
We first evaluated the ability of MACE-OFF23-SC to compute accurate, converged hydration free energies of a series of small organic molecules (Figure 3). For this test, we selected compounds from the FreeSolv database that cover both polar and nonpolar molecules, in addition to several examples known to be challenging for small molecule forcefields [49]. We performed replica exchange molecular dynamics with an ensemble of replicas with $\lambda$ windows spaced in the interval [0,1] to connect the physical end states. To expedite phase space exploration, exchanges of replicas between adjacent $\lambda$ windows were attempted every 1 ps.
For the small systems, we found that 12 replicas enabled sufficient phase-space overlap to converge the free energy simulations (Figure 5). For caffeine and ibuprofen, 20 replicas were required for sufficient exchange probabilities. The replica spacing was hand-tuned from an initial linear spacing to concentrate sampling in the region $0.15\leq\lambda\leq 0.4$, where the curvature of $\frac{\partial H}{\partial\lambda}$ is greatest. For each system, convergence of the free energy as a function of time was confirmed (Figure 4).
Figure 5: Transition probability matrix between 12 replicas from REMD simulation of methanol in water.
Across the range of molecules studied here, we find that free energy predictions computed using MACE-OFF23-SC match or, in many cases, outperform widely-used classical forcefields, with approximately a factor of 2 improvement in MAE and RMSE and improved correlation with experiment (Table 1).
MAE RMSE $r^{2}$
MACE-OFF23-SC 0.53 0.60 0.99
GAFF 2.04 2.59 0.74
OpenFF 2.1 1.79 2.24 0.81
@@ II.2.1 Organic molecule benchmark
For example, compared to OpenFF and GAFF, we find that the MACE-OFF23-SC prediction for methanol falls within the experimental error, whilst the classical forcefields underpredict the magnitude of the hydration free energy. This trend has previously been studied in the context of empirical forcefields and has been attributed to the assignment of partial charges on aliphatic hydroxyl groups [50]. Since our approach is not dependent upon the coarse graining of chemical space through atom-typing, our approach is not susceptible to the same inaccuracies.
To further evaluate the performance of our method, we computed the hydration free energy of several larger molecules, including ibuprofen and caffeine. While still small and rigid compared to a typical therapeutic small-molecule, they are closer to drug-like chemical space than the small organic molecules studied thus far. Due to the larger excluded volume, 20 replicas were required for sufficient phase space overlap, however our method converged quickly to within experimental error. This is particularly encouraging given that, for caffeine, both OpenFF and GAFF predictions have MAEs of around 5 kcal/mol, suggesting that this approach might scale to larger, conformationally flexible drug-like molecules at a similar level of accuracy.
@@ III Conclusion
We have introduced a modified version of the MACE-OFF23 general organic machine learned forcefield that enables, for the first time, theoretically rigorous condensed phase alchemical free energy simulations with first principles based machine learned forcefields. A dual approach of softcore repulsion and $\lambda$-dependent two-body interactions enable numerically well-behaved simulations of weakly interacting atoms, directly analogous to softcore formulations of classical forcefields. Our method enables access to quantities directly relevant to drug discovery at ab initio quality.
We tested the utility of our approach by calculating hydration free energies for a series of small organic molecules. Our method achieves sub-chemical accuracy for a range of polar and nonpolar compounds, improving on state-of-the-art empirical forcefield errors by a factor of two. Equally importantly, we demonstrate that convergence of the free energy is fully feasible within reasonable simulation time, even for the challenging case of disappearing an entire molecule.
We envisage soft-core-equipped machine-learning potentials becoming an increasingly exploited avenue for accurate and efficient calculation of free energies. As well as being an important quantity in forcefield benchmarking, free energies represent key optimisation quantities in drug discovery, where they can give direct access to protein–ligand binding affinities. Hence ab initio quality computational methods are an important tool when rigorous and accurate predictions are required. We further expect that these methods will form part of the benchmarking process for new machine learned forcefields, particularly those incorporating explicit treatment of long range interactions.
@@ IV Methods
IV.1 MACE-OFF23-SC fitting
In previous work, it was found that the use of equivariant features ($L_{max}>0$) in the tensor product significantly increased the accuracy of intermolecular force predictions compared to the invariant MACE-OFF23 model. All simulations in this work were performed using a model with $L_{max}=1$ and 64 channels, which we found to adequately balance computational performance and accuracy.
IV.2 Equilibrium free energy perturbation
Equilibrium free energy calculations were performed using OpenMM, and made use of modified versions of the openmmtools, openmm-ml and openmm-torch libraries to perform Hamiltonian replica exchange and interface MACE with OpenMM [8, 51, 52, 53, 54, 55]. This approach enabled interpolation between the coupled and decoupled endstates via a $\lambda$ dependent Hamiltonian.https://github.com/jharrymoore/mace-md
Since MACE predicts total intermolecular interactions without distinguishing between repulsive exchange and attractive Coulomb and dispersion terms, solvent-solute interactions are switched off using a single transformation, parameterised by the alchemical parameter $\lambda$. We adopt a direct decoupling approach in which the organic molecule is decoupled from a box of water, whilst all other interactions are unmodified. This results in a decoupled end state equivalent to a separated water box and small molecule in vacuum.
Initial solvated structures were generated with pdbfixer, using a rhombic dodecahedral box and a padding distance of 2.0 nm, as required by the 10 Å receptive field of MACE. Structures were energy minimised using the L-BFGS algorithm, and subsequently equilibrated in the NPT ensemble for 50 ps. The last frame was used to seed the replica exchange calculations.
The phase space perturbation was spanned by 12 replicas, with the following spacing [0.0, 0.12, 0.16, 0.20, 0.2425, 0.28, 0.4, 0.52, 0.64, 0.76, 0.88, 1.0], resulting in at least tridiagonal overlap in all cases. Hamiltonian replica exchange between all replicas were attempted every 1000 steps of dynamics (1 ps). It was found that 1 ns of sampling was sufficient to converge free energy estimates in all cases.
Replica exchange calculations were performed using ReplicaExchangeSampler functionality from openmmtools. Each simulation was performed on a single node, comprising 4 NVIDIA A100 GPUs. Replicas were assigned a single MPI rank, and 3 MPI ranks were mapped to each GPU.
@@ IV.3 Empirical forcefield solvation free energies
GAFF solvation free energies and standard errors were taken from the FreeSolv Dataset [49]. OpenFF values were calculated using OpenFF 2.1.0 via the Absolv code, using the default nonequilbrium switching protocolhttps://github.com/SimonBoothroyd/absolv/tree/main.
@@ Acknowledgements
JHM acknowledges support from an AstraZeneca Non-Clinical PhD studentship, and thanks Prof. Ola Engkvist, Dr Marco Klähn and Dr Graeme Robb for their helpful discussions. DJC acknowledges support from a UKRI Future Leaders Fellowship (grant MR/T019654/1). We acknowledge computational resources provided by the Cambridge Service for Data-Driven Discovery (CSD3), the AstraZeneca Scientific Computing Platform (SCP) and the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1
@@
GC has equity stake in Symmetric Group LLP and Ångstrom AI Inc., companies engaged in the application of machine learning to material and molecular simulation.

### v3

@@ II.2.1 Hydration free energies
Figure 5: Transition probability matrix between 16 replicas from REMD simulation of phenol in water.
Across the range of molecules studied here, we find that free energy predictions computed using MACE-OFF24-SC match or, in many cases, outperform widely-used classical forcefields, with a consistent improvement in MAE and RMSE compared to the classical forcefields (Table 1).
MAE RMSE
MACE-OFF24-SC 0.69 0.80
GAFF 1.09 1.24
OpenFF 2.1 0.98 1.15
@@
For example, compared to OpenFF 2.1 [9] and GAFF2, we find that the MACE-OFF24-SC prediction for methanol falls within the experimental error, whilst the classical forcefields underpredict the magnitude of the hydration free energy. This trend has previously been studied in the context of empirical forcefields and has been attributed to the assignment of partial charges on aliphatic hydroxyl groups [55, 56]. Since our approach is not dependent upon the coarse graining of chemical space through atom-typing, our approach is not susceptible to the same inaccuracies.
Notwithstanding atom type specific parametrisation issues, it is encouraging to find that MACE-OFF24-SC is capable of exceeding the accuracy of classical forcefields, especially considering that the nonbonded parameters of both OpenFF 2.1 and GAFF2 have been parametrised to reproduce condensed phase properties of similar small molecules, whereas MACE-OFF24-SC must learn these interactions directly from the quantum mechanical training data. It is also worth noting that all forcefields are approaching the average experimental error of the dataset, around 0.6 kcal/mol, making further quantitative comparison challenging.
II.2.2 Solvation free energies in a non-aqueous solvent
We have shown that MACE-OFF24-SC is capable of highly accurate solvation free energy predictions across a range of small organic molecules. However, whilst these serve as a useful and important benchmark, often the more relevant quantity in drug discovery is the distribution coefficient, which additionally requires the calculation of a solvation free energy in a nonpolar solvent. This probes the ability of the model to describe accurately the intermolecular interactions between the solute and the nonpolar organic solvent, in this case octanol. Although the greater conformational flexibility of the solvent makes convergence of the octanol leg of the thermodynamic cycle slightly more challenging, these calculations were still performed in reasonable computational time, requiring at most 48 hours of walltime (see section II.2.4).
In this section, we evaluate the accuracy of MACE-OFF24-SC predictions on several octanol solvation free energies extracted from the MNSol database [57]. Whilst MNSol does not provide experimental uncertainties, we adopt the best practice developed by the FreeSolv authors and assume an experimental error of 0.6 kcal/mol. With this threshold, all MACE-OFF24-SC predictions fall within the experimental error (Figure 6 and Section S2), indicating that the model is indeed capable of successfully predicting these interactions. Here, we compare to the GAFF2 forcefield equipped with the recently published ABCG2 charge model instead of the normal AM1-BCC, which has been tuned to improve the accuracy of hydration and transfer free energy calculations compared to the standard charge model [56]. Hence, this test represents an extremely stringent comparison for MACE-OFF24-SC, since the empirical forcefield has been directly optimised for the task.
Similar to the hydration free energy results, we find MACE-OFF24-SC has moderately lower prediction error compared to GAFF2/ABCG2, however both values fall below the presumed experimental error, again making further quantitative comparison difficult.
Figure 6: Solvation free energies of a series of small organic molecules in octanol. The shaded region represents a 0.6 kcal/mol deviation.
II.2.3 LogP calculations for drug-like molecules
In order to probe the boundaries of the model’s capabilities, we performed full solvation free energy calculations, in both water and octanol solvents, on a selection of drug-like compounds from the CHEMBL database for which an experimental logP value is available via retention time-based UPLC/MS chromatographic analysis. All compounds satisfy Lipinski’s Rule of 5 criteria [58] and are representative of the chemical complexity and flexibility of molecules typical of a medicinal chemistry campaign. Structures of the compounds are shown in Section S9.
Using the same alchemical free energy methodology as for the earlier calculations, we calculated logP as the difference between the water and octanol free energy legs:
$\log P=\frac{\Delta G_{\text{hyd}}-\Delta G_{\text{oct}}}{2.303RT}$ (7)
Figure 7: Experimental logP for CHEMBL compounds compared to OpenFF 2.1 and MACE-OFF24-SC. *Experimental value recorded as $>5.4$, so quantitative comparison is not possible.
Figure 7 compares both MACE-OFF24-SC and OpenFF 2.1 with experimental measurements. Here, for the first time we see MACE-OFF24-SC significantly outperforming the classical forcefield, with an RMSE of 0.45 log units, compared to 4.02 for OpenFF 2.1. This result is extremely encouraging, suggesting that the model is capable of transferring nonbonded interactions learned from the training set to highly functionalised, conformationally complex molecules. This result validates our hypothesis that machine learning forcefields, which have learned these interactions from first principles, may show superior transferability to drug-like chemical space, compared to the explicit parametrisation approach used to fit empirical forcefields.
II.2.4 Computational performance
Free energy calculations, even with classical forcefields, are known to be computationally expensive. Hydration free energy calculations ameliorate the sampling problem for the purposes of benchmarking, since the phase space perturbations are small and expected to converge rapidly compared to protein-ligand binding free energies. Running free energy calculations with machine learned potentials raises the issue of whether converged results can be obtained within reasonable wall time.
Here, all calculations were performed on single nodes, containing either 8 NVIDIA A100, or 8 NVIDIA L40S, GPUs. For both machine types, we achieved aggregated sampling of around 8 ns/day across 16 replicas, corresponding to approximately 500 replica exchange iterations of 1 ps each per day. Most hydration free energies converged within this time, whilst some were extended to run for 1 ns per replica, requiring 48 hours of wall time. MNSol octanol solvation free energies were all run for 1 ns per replica to account for the additional conformational sampling requirement.
Owing to the size of the solutes in the logP calculations, these calculations took longer to converge, with the average hydration and octanol leg requiring 4 and 7 days of walltime, respectively.
@@ III Conclusion
We have introduced a modified version of the MACE-OFF24 transferable organic machine learned forcefield that enables, theoretically rigorous condensed phase alchemical free energy simulations with first principles based machine learned forcefields. A dual approach of softcore repulsion and $\lambda$-dependent two-body interactions enable numerically well-behaved simulations of weakly interacting atoms, directly analogous to softcore formulations of classical forcefields. Our method enables access to quantities directly relevant to drug discovery at ab initio quality.
We tested the accuracy of our approach by calculating free energies of solvation, in water and octanol solvents, for a range of small organic molecules from the FreeSolv and MNSol databases [53, 57]. Our approach achieves excellent accuracy on a wide range of functional groups, improving on state-of-the-art empirical forcefield accuracy and approaching the experimental errors of the datasets.
As a final stringent test of the accuracy of the model, we computed logP values for a series of compounds from the CHEMBL database. These compounds are drug-like in terms of size, complexity and functionality, are expected to be neutral at physiological pH and have experimental logP values. On this benchmark, we significantly outperform OpenFF 2.1, a state-of-the-art classical forcefield, decreasing the RMSE by an order of magnitude. This performance is highly encouraging for future applications of machine learning forcefields in drug discovery, and represents a significant step towards replacing labour and resource intensive chemical synthesis and experiment with fast and accurate molecular simulation.
Equally importantly, we demonstrate that convergence of the free energy is feasible within reasonable simulation time, even for absolute solvation free energy calculations in octanol, which present additional sampling challenges compared to aqueous free energy calculations. Since this work was performed, optimised computational kernels such as NVIDIA’s cuEquivariance have been releasedhttps://github.com/NVIDIA/cuEquivariance, which we expect to significantly improve the throughput of these calculations.
As has been previously noted for MACE-OFF models, the lack of an explicit long-range contribution limits the chemical space in which the model is expected to be accurate. This has been addressed by recent transferable MLPs including AIMNet2 [27] and FeNNix-Bio1 [46] which include an explicit Coulomb term. The latter employs the alchemical methods developed in this work and uses solvation and protein-ligand binding free energies to benchmark the accuracy of the model for condensed phase systems.
We believe softcore-equipped machine learning potentials will become an increasingly exploited avenue for accurate and efficient calculation of free energies. As well as being an important quantity in forcefield benchmarking, free energies represent key optimisation quantities in drug discovery, where they can give direct access to protein–ligand binding affinities. Hence accurate and efficient computational methods are an important tool when rigorous and accurate predictions are required.
@@ IV Methods
IV.1 MACE-OFF24-SC training
In previous work, it was found that the use of equivariant features ($L_{max}>0$) in the tensor product significantly increased the accuracy of intermolecular force predictions compared to the invariant MACE-OFF23 model [25]. All simulations in this work were performed using a model with $L_{max}=1$, 128 channels, and a layer-wise cutoff of 6 Å. These hyperparameters were adopted from the MACE-OFF24(M) model, given its ability to accurately describe the properties of condensed phase systems [25].
MACE-OFF24-SC was trained on the same neutral-only filtering of the SPICE2 dataset that we previously reported, which included, crucially, the newly added protein-ligand and solvated pubchem datasets, which provided extensive sampling of the intermolecular interactions required for accurate liquid phase free energy calculations [59].
IV.2 Equilibrium free energy calculations
Equilibrium free energy calculations were performed using OpenMM, and made use of modified versions of the openmmtools, openmm-ml and openmm-torch libraries to perform Hamiltonian replica exchange and interface MACE with OpenMM [8, 60, 61, 62, 63, 64]. This approach enabled interpolation between the coupled and decoupled endstates via a $\lambda$-dependent Hamiltonian.https://github.com/jharrymoore/mace-md
Since MACE predicts total intermolecular interactions without distinguishing between repulsive exchange and attractive Coulomb and dispersion terms, solvent-solute interactions are switched off using a single transformation, parameterised by the alchemical parameter $\lambda$. We adopt a direct decoupling approach in which the organic molecule is decoupled from the surrounding solvent, whilst all other interactions are unmodified. This results in a decoupled end state equivalent to a separated solvent box and small molecule in vacuum.
Initial solvated structures were generated with pdbfixer, using a rhombic dodecahedral box and a padding of 1.3 nm from solute to box edge, ensuring the minimum image convention is correctly applied for the 12 Å receptive field of MACE. Structures were energy minimised using the L-BFGS algorithm, and subsequently equilibrated in the NPT ensemble for 50 ps. The last frame was used to seed the replica exchange calculations.
Replicas were then propagated in the NPT ensemble under Langevin dynamics with an integration timestep of 1 fs and a friction coefficient of 1/ps. Pressure was maintained at 1 atm with a MonteCarlo barostat as implemented in OpenMM.
The phase space perturbation was spanned in all cases by 16 replicas, which resulted in at least tridiagonal overlap in all cases. Hamiltonian replica exchange between all replicas were attempted every 1000 steps of dynamics (1 ps). It was found that 1 ns of sampling per replica was sufficient to converge the free energy estimates in all cases. A full sample input file is provided in Section S7
Replica exchange calculations were performed using ReplicaExchangeSampler class from openmmtools. Replicas were assigned a single MPI rank, and two MPI ranks were mapped to each GPU, enabling full utilisation of the GPU’s compute bandwidth, thus slightly increasing throughput compared to running the replicas in multiple waves.
Automated equilibration detection and trajectory subsampling were performed following Chodera et al. as implemented in pymbar to ensure only decorrelated samples were used to solve the MBAR equations [65, 41].
@@ IV.3 Empirical forcefield solvation free energies
GAFF solvation free energies and standard errors were taken from the FreeSolv Dataset [53]. OpenFF values were calculated using OpenFF 2.1.0 via the Absolv code, using an equilibrium free energy protocolhttps://github.com/SimonBoothroyd/absolv/. An example script can be found in Section S8.
@@ Acknowledgements
JHM acknowledges support from an AstraZeneca Non-Clinical PhD studentship, and thanks Prof. Ola Engkvist, Dr Marco Klähn and Dr Graeme Robb for their helpful discussions. DJC acknowledges support from a UKRI Future Leaders Fellowship (grant MR/T019654/1). We acknowledge computational resources provided by the Cambridge Service for Data-Driven Discovery (CSD3), the AstraZeneca Scientific Computing Platform (SCP), the UK national high-performance computing service, ARCHER2, for which access was obtained via the UKCP consortium and the EPSRC grant ref EP/P022561/1, and Ångström AI.
@@
GC has equity stake in Symmetric Group LLP. GC and JHM have equity stake in Ångström AI Inc. Both companies are engaged in the application of machine learning to material and molecular simulation.
Supporting Information for:
Computing solvation free energies of small molecules with first principles accuracy
J. Harry Moore Daniel J. Cole
Gábor Csányi
Appendix S1 Training and test set energy and force errors
Table S2: Route mean squared errors (RMSE) for training, validation and test set configurations for MACE-OFF24-SC. See previous work for details of the datasets [25]. Energy errors are in meV/atom, and force errors in meV/Å.
Dataset type E RMSE F RMSE Rel. F RMSE %
Training 1.2 28.7 3.96
Validation 1.2 28.9 3.70
DES370K Dimers 2.1 15.2 3.23
DES370K Monomers 2.5 16.4 2.54
Dipeptides 2.5 22.9 3.25
PubChem 2.8 33.7 4.36
QMugs 2.3 35.5 3.53
Solvated Amino Acids 1.3 33.4 2.73
Water 1.9 23.6 3.72
Appendix S2 Octanol solvation free energy results
Molecule Experiment Experiment Error MACE MACE Error GAFF2/ABCG2
methanol -3.87 0.6 -3.87 0.09 -3.223
phthalimide -11.18 0.6 -10.25 0.21 -10.945
methane 0.51 0.6 0.26 0.15 0.64
thiophene -3.89 0.6 -3.05 0.11 -3.807
benzonitrile -6.09 0.6 -6.56 0.15 -6.169
caffeine -12.54 0.6 -11.73 0.17 -15.708
bromomethane -2.43 0.6 -1.58 0.2 -2.264
o-cresol -8.49 0.6 -8.95 0.16 -8.58
benzene -3.72 0.6 -3.7 0.11 -3.121
acetone -3.15 0.6 -3.26 0.19 -2.935
Table S3: Solvation free energies in octanol for a series of organic solutes. GAFF2-ABCG2 values are taken from the literature. [56]
Appendix S3 FreeSolv hydration free energy results
Compound Experimental MACE GAFF
mobley_2789243 -4.01 -2.9 -2.77
mobley_1244778 -5.48 -6.44 -4.34
mobley_6091882 1.28 0.33 2.33
mobley_4291494 -4.39 -3.77 -3.87
mobley_8966374 -5.51 -4.65 -6.16
mobley_1881249 -3.52 -4.64 -3.17
mobley_4884177 -4.59 -5.78 -2.96
mobley_6102880 -2.87 -3.93 -2.81
mobley_7039935 -7.67 -7.2 -7.14
mobley_9029594 -0.22 -0.9 0.88
mobley_2782339 -9.61 -8.49 -11.82
mobley_7326982 -4.74 -5.44 -3.43
mobley_8764620 -6.75 -7.82 -5.58
mobley_6353617 -8.26 -7.52 -7.84
mobley_8899867 -5.49 -5.76 -4.29
mobley_6303022 -4.73 -5.62 -3.19
mobley_6794076 -7.62 -7.3 -7.57
mobley_4188615 -9.29 -8.49 -8.11
mobley_1963873 -10.0 -10.43 -8.28
mobley_4639255 -9.3 -10.63 -7.27
mobley_8124669 -8.21 -7.27 -6.85
mobley_7608462 -1.12 -0.48 -0.47
mobley_8578590 -9.28 -10.88 -8.24
mobley_6988468 -4.33 -5.16 -3.31
mobley_6861308 -9.13 -10.39 -11.81
mobley_5220185 -5.45 -5.14 -5.81
mobley_1352110 -3.64 -3.81 -4.57
mobley_2316618 -14.21 -14.34 -13.6
mobley_7010316 -5.46 -6.35 -4.18
mobley_6739648 -5.06 -5.94 -4.27
mobley_1903702 -4.62 -5.22 -3.15
mobley_2049967 -1.5 -2.04 0.39
mobley_9055303 2.0 2.15 2.45
mobley_1636752 -5.1 -5.09 -3.49
mobley_2008055 1.83 1.92 2.46
mobley_2310185 -5.0 -4.98 -3.39
mobley_20524 -6.6 -6.36 -5.71
Table S4: Results for MACE and GAFF on the freesolv subset compared to experiment
Appendix S4 Switching points
Dimer pair Switching point / Å Dimer pair Switching point / Å
Br2 2.1499996 FO 1.2799997
BrCl 2.0199995 FP 1.5199996
BrF 1.6749996 FS 1.5249996
BrI 2.3149996 H2 0.6749997
BrN 1.6399996 HBr 1.3299996
BrO 1.6299996 HCl 1.2049997
BrP 2.0249996 HF 0.87999976
BrS 2.0049996 HI 1.4949995
C2 1.2149997 HN 0.9749997
CBr 1.7049996 HO 0.9199998
CCl 1.5549996 HP 1.3099995
CF 1.2149997 HS 1.2499996
CH 1.0349996 I2 2.480009
CI 1.8849996 IN 1.8049996
CN 1.1349998 IO 1.7749996
CO 1.0999999 IP 2.1949997
CP 1.4949996 IS 2.1649995
CS 1.4749997 N2 1.0699998
Cl2 1.8849995 NO 1.1149998
ClF 1.5499996 NP 1.4349997
ClI 2.1799996 NS 1.4399997
ClN 1.4899997 O2 1.1649997
ClO 1.4899997 OP 1.4299997
ClP 1.8899995 OS 1.4249997
ClS 1.8749995 P2 1.8049997
F2 1.3249997 PS 1.8149996
FI 1.8149996 S2 1.8049996
FN 1.2399998
Appendix S5 Example of nonequilibrium switching
Figure S8: Example $\frac{dH}{d\lambda}$ plot for the solvation of methane in water over a 35 ps nonequilibrium switching trajectory.
Appendix S6 Free energy convergence plots
Figure S9: Convergence of hydration free energy calculations from FreeSolv benchmark.
Appendix S7 Sample replica exchange input file
1#!/bin/bash
2mpirun -np 8 mace-md -f /path/to/equilibrated/structure.pdb \
3 --resname "LIG"\
4 --output_dir /path/to/output/dir \
5 --pressure 1.0 \
6 --decouple \
7 --restart \
8 --dtype float32 \
9 --log_level INFO \
10 --run_type "repex" \
11 --system_type "pure" \
12 --interval 1 \
13 --steps 1000 \
14 --lambda_schedule="[0. , 0.08, 0.1, 0.13, 0.15, 0.17, 0.20, 0.25, 0.2825, 0.385 , 0.4875, 0.59 , 0.6925, 0.795 , 0.8975, 1.]"\
15 --model_path "/path/to/mace/model" \
16 --steps_per_iter 1000 \
Listing 1: Script to run 1 ns of replica exchange simulation via the MACE-MD code
Appendix S8 Example Absolv run script
1import pathlib
2import sys
3import openmm.unit
4import openff.toolkit
5from absolv import config, runner
6import absolv
7
8smiles = sys.argv[1]
9temperature=298.15 * openmm.unit.kelvin
10pressure=1.0 * openmm.unit.atmosphere
11
12system=absolv.config.System(
13 solutes={smiles: 1}, solvent_a=None, solvent_b={"O": 1500}
14)
15
16alchemical_protocol_a=absolv.config.EquilibriumProtocol(
17 lambda_sterics=[
18 1.00, 1.00, 1.00, 1.00, 1.00, 0.95, 0.90, 0.80, 0.70, 0.60, 0.50, 0.40,
19 0.35, 0.30, 0.25, 0.20, 0.15, 0.10, 0.05, 0.00,
20 ],
21 lambda_electrostatics=[
22 1.00, 0.75, 0.50, 0.25, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00,
23 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00,
24 ]
25)
26alchemical_protocol_b=absolv.config.EquilibriumProtocol(
27 lambda_sterics=[
28 1.00, 1.00, 1.00, 1.00, 1.00, 0.95, 0.90, 0.80, 0.70, 0.60, 0.50, 0.40,
29 0.35, 0.30, 0.25, 0.20, 0.15, 0.10, 0.05, 0.00,
30 ],
31 lambda_electrostatics=[
32 1.00, 0.75, 0.50, 0.25, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00,
33 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00, 0.00,
34 ]
35)
36
37config = absolv.config.Config(
38 temperature=temperature,
39 pressure=pressure,
40 alchemical_protocol_a=alchemical_protocol_a,
41 alchemical_protocol_b=alchemical_protocol_b,
42)
43
44force_field = openff.toolkit.ForceField("openff-2.1.0.offxml")
45prepared_system_a, prepared_system_b = absolv.runner.setup(system, config, force_field)
46
47result = absolv.runner.run_eq(
48 config, prepared_system_a, prepared_system_b, "CUDA"
49)
50
51path = pathlib.Path("results")
52(path / "output.txt").write_text(result.model_dump_json(indent=2))
Appendix S9 CHEMBL compounds
Figure S10: Structures of the CHEMBL compounds used for the logP calculations.